OpenStudy (anonymous):

pumping 1 𝑑mˆ3 air per second into a balloon. How fast increasing radius when the diameter is 20 𝑐𝑚?

OpenStudy (campbell_st):

well the volume of the balloon is \[V = \frac{4}{3} \pi r^3\] so find \[\frac{dV}{dr}\] when you find the derivative, then find the reciprocal you will need \[\frac{dr}{dV} \] later you are given \[\frac{dV}{dt} = 1\] and you need to find \[\frac{dr}{dt}\] so you can use the chain rule \[\frac{dr}{dt} = \frac{dr}{dV} \times \frac{dV}{dt}\] hope it makes sense

OpenStudy (campbell_st):

when you get \[\frac{dr}{dt}\] when the diameter is 20, the radius is 10, so substitute r = 10 and evaluate